# The time scale of redundancy between prosody and linguistic context

### v2

@@ 3.2 Manipulating Context Length
Thus, for instance, $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{0,0}{\leftrightarrow}}}$ corresponds to the word ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}_{t}}}$ by itself, and $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{3,6}{\leftrightarrow}}}$ corresponds to the word ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}_{t}}}$ with 3 words in its past context and 6 words in its future context (see Fig. 2).
Given this definition, we can then explore the time scale we are interested in by estimating the mutual information ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}},{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$ while varying $n$ and $m$. This amounts to estimating $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}})$ and $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$.
@@
The unconditional entropy does not depend on context; thus, to estimate it, we only need to compute a relatively simple prior over the domain of each prosodic feature. On the other hand, estimating the conditional entropy requires computing a family of conditional distributions for each prosodic feature, with one distribution for each $n,m$ combination.
In other words, we need a model $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$ that works for any $n,m$ pair. We elaborate on this model in Section § 4.2.
@@
Importantly, the ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}};{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$ is a monotonically increasing function of both $m$ and $n$; larger linguistic contexts must contain at least as much (but maybe more) information about prosody than smaller contexts. [^4] We note that—while the MI is monotonically increasing in theory—it is not necessarily the case that this underlying monotonicity will be reflected by our estimation methods. However, at some value of $m$ and some value of $n$, the MI might reach a plateau. We consider the “time scale” of the redundancy between prosody and linguistic context to be the value from which increasing context does not significantly increase the MI.
@@ 4.1 Dataset
Our data-extraction process follows Wolf et al.’s (2023) proposed and publicly available pipeline (see their paper for more details).
@@
We use the LibriTTS spoken language corpus  [^5] The LibriTTS corpus is licensed under the Creative Commons Attribution 4.0 International License (CC BY 4.0) Zen et al. (2019), [^6] This dataset is derived from LibriSpeech audiobooks corpus Panayotov et al. (2015), which is itself derived from LibriVox Kearns (2014). which contains public domain audiobook materials (audio and text) recorded by volunteer narrators. This dataset contains 585 hours of English speech data at a 24kHz sampling rate, and includes recordings from 2,456 speakers reading aloud books which are paired with the corresponding transcripts. We filtered out texts from LibriTTS that contained less than three words (such as book and chapter titles) and we eliminated punctuation marks, since these can be very informative regarding prosody, and are not explicitly present in spoken communication.
@@
The procedure for extracting prosodic features starts with aligning the audio and text using Montreal Forced Aligner (MFA) McAuliffe et al. (2017).
Given this alignment, both the duration and pause of each word can be easily computed from the words’ offset and onset times. [^7] We note that about $89.4\%$ of the words in our dataset have a pause of 0 seconds. Duration was then normalized relative to the number of syllables in the word, to reflect duration per syllable.
@@
In short, this prominence value reflects the steepest time-frequency variation in a signal combining duration, energy and $f_{0}$. Apart from the absolute average prominence value per each word, we also computed their relative prominence, subtracting the mean prominence value of the three preceding words from the current word’s value; this emphasizes local changes in prominence.
Splitting the Data to Train, Validation and Test
The dataset was divided into separate train, validation and test sets, using splits of the dataset provided by Talman et al. (2019). For training, we used a data split (termed train-360) containing 904 speakers, 11,262 sentences and 2,076,289 words. For validation, we used a data split (termed dev) containing 40 speakers, 5,726 sentences and 99,200 words. We had access to a data split (termed test) containing 39 speakers, 4,821 sentences and 90,063 words, as well as to another data split (termed train-100) containing 247 speakers, 33,041 sentences and 570,592 words. We therefore used the latter split (train-100) as the test set, except for prominence (absolute and relative) where we added this split to the test split, which led to more stable results.
@@ 4.2 Estimating the Cross-Entropies
To model this conditional distribution, we again follow Wolf et al. in finetuning a language model ($\mathrm{LM}$), with an added linear layer on top, to predict the parameters of a conditional distribution over ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}$. Unlike Wolf et al., however, we limit our models input to include only ${\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}}$ instead of the entire ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}$. We assume the conditional distribution over prosody to follow a parametric distribution $\mathcal{Z}$, and use a $\mathrm{LM}$ to predict this distribution’s parameters $\widehat{\boldsymbol{\phi}}$: [^8] We evaluate models with Gaussian, Gamma and Laplace distributions, choosing the distribution that leads to the lowest cross-entropy on a validation set. Parameters $\boldsymbol{\phi}$ are, e.g., the mean and standard deviation for a Gaussian.
@@
We finetune this model by minimizing its cross-entropy on a training set $\mathcal{D}_{\mathrm{trn}}$; which amounts to minimizing the right-hand side of Eq. 2. As the cross-entropy is an upper bound on the entropy, the lower its value (and consequently the better our model) the tighter the estimate we get for the entropy. Notably, we estimate all $n,m$ combinations using a single finetuned $\mathrm{LM}$. During training, we sampled inputs of varying lengths, spanning 1 to 10 words. [^9] These inputs were obtained by first sampling an item from the dataset, and then randomly cutting it into the desired length (between 1-10 words). For each sample, the model then predicts the prosody of each of the words in this span in parallel; in a 7-word span, thus, the first word’s prosody is predicted as $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{0,6}{\leftrightarrow}}}})$ and the 5th word’s prosody is predicted as $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{4,2}{\leftrightarrow}}}})$.101010 For each prosodic feature, we tested several models, namely BERT, BERT-large and RoBERTa-large, and selected the model that gave the best results. In most cases, this was BERT-large, except for pauses and pitch where it was BERT. An early stopping criterion was applied, such that if the loss did not decrease for 3 epochs the model stopped training.
@@
Figure 3: For each of the 6 tested prosodic features, two plots are presented. The upper plots are similar to Fig. 1, and display the redundancy, quantified as mutual information between a prosodic feature at a given word $\mathrm{p}_{t}$and the linguistic context $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}$, which includes the word itself, $n$ words before, and $m$ words after it. The lower plots display just the first column (corresponding to linguistic context which includes the word itself and a gradually increasing number of past words, red curve) and the first row (corresponding to lingusitic context which includes the word itself and a gradually increasing number of past words, blue curve) in the upper plots. Wherever numbers and units are not displayed, they correspond to the units displayed for Absolute Prominence. The MI values correspond to the mean across all train data. Error bars correspond to standard errors of the mean. Dashed horizontal line corresponds to the MI with the longest available context. See App. B for the distributions of these features from our dataset.

### v3

@@ 3.2 Manipulating Context Length
Thus, for instance, $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{0,0}{\leftrightarrow}}}$ corresponds to the word ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}_{t}}}$ by itself, and $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{5,3}{\leftrightarrow}}}$ corresponds to the word ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}_{t}}}$ with 5 words in its past context and 3 words in its future context (see Fig. 2).
@@
Given this definition, we can then explore the time scale we are interested in by estimating the mutual information ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}},{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$ while varying $n$ and $m$. This amounts to estimating $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}})$ and $\mathrm{H}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$. The unconditional entropy does not depend on context; thus, to estimate it, we only need to compute a distribution over the domain of each prosodic feature. On the other hand, estimating the conditional entropy requires computing a family of conditional distributions for each prosodic feature, with one distribution for each $n,m$ combination. In other words, we need a model $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$ that works for any $n,m$ pair. We elaborate on this model in section § 4.2.
@@
Importantly, the ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}};{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$ is a monotonically increasing function of both $m$ and $n$; larger contexts must contain at least as much (but maybe more) information about prosody than smaller contexts. [^3] We note that—while the MI is monotonically increasing in theory—it is not necessarily the case that this underlying monotonicity will be reflected by our estimation methods. However, at some value of $m$ and some value of $n$, the MI might reach a plateau. We consider the time scale of the redundancy between prosody and linguistic context to be the value from which increasing context length does not significantly increase the MI.
@@ 4.1 Dataset
Our data-extraction process uses Wolf et al.’s (2023) proposed and publicly available pipeline (see their paper for more details).
@@
We use the LibriTTS spoken language corpus Zen et al. (2019), [^4] This dataset is derived from LibriSpeech audiobooks corpus Panayotov et al. (2015), which is itself derived from LibriVox Kearns (2014). which contains public domain audiobook materials (audio and text) recorded by volunteer narrators. This dataset contains 585 hours of English speech data at a 24kHz sampling rate, and includes recordings from 2,456 speakers reading aloud books which are paired with the corresponding transcripts. We filtered out texts from LibriTTS that contained less than three words (such as book and chapter titles) and we eliminated punctuation marks, since these can be very informative regarding prosody, and are not explicitly present in spoken communication.
@@
For extracting prosodic features we began by aligning the audio and text using the Montreal Forced Aligner (MFA; McAuliffe et al., 2017). With the alignment in place, both the duration and pause of each word can be easily computed from the words’ offset and onset times. Notably, $89.4\%$ of the words in our dataset have a pause of 0 seconds. Duration was normalized by the number of syllables (from CELEX; Baayen et al., 1996) to obtain duration per syllable, reducing variability across word identities.
@@
Prominence was computed using a composite acoustic measure from Suni et al. (2017), which integrates time-frequency variation across duration, energy, and $f_{0}$. This measure has been validated in prior work as capturing prosodic prominence Suni et al. (2017); Talman et al. (2019); Terken and Hermes (2000). In addition to the absolute prominence values, we also computed relative prominence by subtracting the mean prominence of the preceding three words from the current word’s value, emphasizing local contrast. We used the word-level prominence scores released with this dataset Talman et al. (2019).
Splitting the Data into Train, Validation and Test
The dataset was divided into train, validation and test sets, following standard practice in machine learning. We used splits of the dataset provided by Talman et al. (2019). For training, we used a data split (termed train-360) containing 904 speakers, 11,262 sentences and 2,076,289 words. For validation, we used a data split (termed dev) containing 40 speakers, 5,726 sentences and 99,200 words. For testing, we used two data splits (both non-overlapping with the Train and Validation sets): one (termed train-100) containing 247 speakers, 33,041 sentences and 570,592 words, and the other (termed test) containing 39 speakers, 4,821 sentences and 90,063 words. We used the train-100 split as the main test set, except when analyzing prominence; for those analyses, we combined the two test sets in order to obtain more stable results.
@@ 4.2 Estimating the Cross-Entropies
To model this conditional distribution, we again follow Wolf et al. in finetuning a language model ($\mathrm{LM}$), with an added linear layer on top, to predict the parameters of a conditional distribution over ${{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}$. Using language models to represent text offers greater expressivity than traditional regression-based approaches. Unlike Wolf et al., however, we limit our model’s input to include only ${\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}}$ instead of the entire ${{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}$ (Wolf et al. used the entire available context in the segments from the LibriTTS corpus, which varied in length between 4 and 60 words).
We assume the conditional distribution over prosody follows a parametric distribution $\mathcal{Z}$, and use a $\mathrm{LM}$ to predict this distribution’s parameters $\widehat{\boldsymbol{\phi}}$: [^5] We evaluate models with Gaussian, Gamma and Laplace distributions, choosing the distribution that leads to the lowest cross-entropy on a validation set. Parameters $\boldsymbol{\phi}$ are, e.g., the mean and standard deviation for a Gaussian.
@@
We finetune this model by minimizing its cross-entropy on a training set $\mathcal{D}_{\mathrm{trn}}$; which amounts to minimizing the right-hand side of Eq. 2. As the cross-entropy is an upper bound on the entropy, the lower its value (and consequently the better our model) the tighter the estimate we get for the entropy. Due to compute time considerations, we estimate all $n,m$ combinations using a single finetuned $\mathrm{LM}$: During training, we sampled inputs of varying lengths, spanning 1 to 10 words. These inputs were obtained by first sampling an item from the dataset, and then randomly cutting it into the desired length (between 1-10 words). For each sample, the model then predicts the prosody of each of the words in this span in parallel; in a 7-word span, thus, the first word’s prosody is predicted as $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{0,6}{\leftrightarrow}}}})$ and the 5th word’s prosody is predicted as $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{4,2}{\leftrightarrow}}}})$. Importantly, we made sure that the models saw an equal number fo samples for each $n,m$ combination. For each prosodic feature, we tested several models, namely BERT, BERT-large and RoBERTa-large, and selected the model that gave the best results. In most cases, this was BERT-large, except for pauses and pitch where it was BERT. An early stopping criterion was applied, such that if the loss did not decrease for 3 epochs the model stopped training.
@@
Figure 3: For each of the 6 tested prosodic features, two plots are presented. The upper plots are similar to Fig. 1, and display the redundancy, quantified as mutual information between a prosodic feature at a given word $\mathrm{p}_{t}$ and the linguistic context $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}$, which includes the word itself and a varying length context, $n$ words before, and $m$ words after it. The lower plots display either: (a) in red, just the first column from the upper plots (corresponding to past linguistic context, which includes the word itself and a gradually increasing number of previous words, red curve) and (b), in blue, the first row from the upper plots (corresponding to future linguistic context, which includes the word itself and a gradually increasing number of future words, blue curve). Details displayed only for Pitch or Duration (labels, numbers, legend) are the same for all features. The MI values correspond to the mean across all test data. Error bars correspond to standard errors of the mean. Dashed horizontal gray lines corresponds to the MI with the longest available context, taken from Wolf et al. (2023) (between 6 and 60 words into the past or future). See section App. B for the distributions of these features in our dataset.